# EMORL: Ensemble Multi-Objective Reinforcement Learning for Efficient and Flexible LLM Fine-Tuning

### v2

@@ 5 Experiments
Figure 5: The visualization illustrates the hierarchical grid search process, showing the transition from broad search spaces to more refined spaces, where optimal weight combinations are identified. The red line on the color map indicates the maximum mean reward achieved during the search.
@@ 5.1 Models
We employ T5-base [^1] https://huggingface.co/google-t5/t5-base (220M parameters, Encoder-Decoder architecture) as the pre-trained model in this paper. Pérez-Rosas et al. (2024) demonstrated T5’s effectiveness for understanding and generating text in counseling tasks, making it suitable for our experiments compared with existing baselines. We utilize Self-Critical Sequence Training (SCST) as the RL algorithm, which generates candidate outputs and computes their mean reward as a baseline, thereby encouraging outputs to exceed this baseline performance (Laban et al., 2021). This approach integrates KL-divergence in the loss function to constrain the fine-tuned model from diverging too far from the reference pre-trained model.
We fine-tune models for reflection, empathy, and fluency objectives across $5$ random seeds, with $3$ generation runs per seed, a training batch size of $16$, and up to $10,000$ steps while implementing early stopping at model convergence. We also employ LoRA Hu et al. (2021) to efficiently manage parameter updates by representing them via low-rank matrices and a scaling factor, as illustrated in Equation (4). This approach allows us to load the pre-trained model only once during inference and apply the LoRA parameters to update the pre-trained model toward each objective, significantly reducing the computational burden.
For evaluation, we compared EMORL models against four baselines: T5-base, Uniform Weighted, DynaOpt (Pérez-Rosas et al., 2024), and Model Soups (Wortsman et al., 2022). All experiments were conducted on a Tesla V100 GPU with 32GB memory, 8 CPU cores, and 40GB system memory, with detailed consumption reported in Table 1.
@@ 5.2 Datasets
Counselor reflection generation task is a single-turn task to generate therapist-like reflective statements that accurately paraphrase and validate a client’s expressed thoughts and feelings. It reflects the seeker’s problem and establishes the first step for therapeutic communication. The responses are expected to be reflective, empathic to the problem, and maintain fluency, as depicted in Figure 1.
The PAIR [^2] https://lit.eecs.umich.edu/downloads.html dataset is our primary dataset, split into a ratio of $80\%$, $10\%$, and $10\%$, for training, aggregation, and inference, respectively (Min et al., 2022). It contains $2,544$ single-turn client-counselor exchanges, covering topics ranging from mental health to lifestyle concerns like diet, exercise, and personal development. To assess the models robustly, we also conduct evaluations on the Psych8k [^3] https://huggingface.co/datasets/EmoCareAI/Psych8k dataset, sampling $10\%$ of its $8,187$ conversation pairs for inference. This dataset focuses on mental health interactions, including anxiety, depression, relationship issues, and stress management. It is widely used for training and evaluating LLMs in mental health counseling, and we leverage it here for generating reflective statements.
@@ 5.3 Metrics
We evaluate comprehensive metrics beyond performance, focusing on six key aspects: (1) Diversity-2 measures linguistic diversity; (2) Edit rate quantifies the avoidance of verbatim repetition; (3) Data consumption tracks the cumulative number of training samples processed; (4) Time consumption records the wall-clock time for each training iteration; (5) Scalability assesses the model’s ability and flexibility to incorporate additional objectives; (6) Explainability examines the interpretability of how each objective contributes to the final output.
For performance metrics, we employ specific LLMs to score three objectives on a scale from $0.0$ to $1.0$: (1) Reflection is assessed by the "roberta-base [^4] https://huggingface.co/FacebookAI/roberta-base" with checkpoints from Pérez-Rosas et al. (2024), which evaluates the relevance and contextual appropriateness. (2) Empathy is measured by the "bert-empathy [^5] https://huggingface.co/MoaazZaki/bert-empathy", which gauges emotional resonance and understanding. (3) Fluency is evaluated using "gpt2666https://huggingface.co/openai-community/gpt2" by computing the inverse of perplexity, ensuring linguistic smoothness.
@@ 6 Results Analysis
Hierarchical Grid Visualization Demonstrates Explainable Results.
Figure 5 illustrates the hierarchical grid search process: in iteration $1$ (left subplot), we evaluated $3\times 3\times 3$ weighted combinations for reflection, empathy, and fluency, with color mapping indicating mean reward values. The $2\times 2\times 2$ area with best total performance was then used to refine the search space for subsequent iterations. The right subplot of Figure 5 shows iteration $5$, where the search converged to a more precise and refined space: $(0.75,0.8125)$, $(0.4375,0.5)$ and $(0.0,0.0625)$ for the aggregation weights of reflection, empathy, and fluency models respectively. This progressive refinement allows for the precise identification of the optimal weighted combination.
The visualization reveals significant insights regarding each objective’s contribution to overall performance. The aggregation benefits (orange points) from higher weights of the reflection model. Empathy delivers optimal overall performance with moderate weights, while extreme weights reduce the mean reward. The fluency model demonstrates a negative effect (cyan points) on other objectives when assigned high weights, with lower weights facilitating better integration with the other objectives. This visualization can be readily acquired during aggregation by evaluating weighted combinations on batches of test data. Although results vary across experiments with different sampled models, this approach underscores the interpretability and explainability of EMORL.
EMORL Shows Promise Across Multiple Evaluation Metrics.
$T_{total}=\max_{i\in\{1,2,3\}}\{T_{train}(obj_{i})\}+T_{agg}$ (3)
This parallel process enables over 0.5× faster training compared to single-policy methods. While another meta-policy method, Model Soups, employs gradient-based algorithms (learned soup) that often struggle with local optima, leading to increased consumption. However, EMORL employs hidden-state level aggregation and requires token-by-token generation, which interrupts the sequential generation process of transformers, slightly increasing time consumption. EMORL exhibits greater stability, with variations of only $1,650$ data points and $147.43$ seconds, significantly lower than DynaOpt’s variations of $16,736$ data points and $4,365.64$ seconds. This stability is attributed to EMORL’s consistent single-objective training and uniform optimization resource consumption facilitated by hierarchical grid search. These advantages position EMORL as an efficient and stable fine-tuning approach for multi-objective tasks.
EMORL is a flexible fine-tuning approach that achieves both scalability and explainability. Its scalability is demonstrated when adding new objectives: instead of retraining existing models, EMORL only trains the new single-objective model and optimizes the weights through the aggregation phase. For explainability, EMORL provides clear insights through weighted combinations and hidden-state level aggregation patterns. Figure 5 reflects varying contributions: reflection approximately $0.8$, empathy around $0.5$, and fluency about $0.05$. Adjusting one objective’s weight enhances its performance while affecting other objectives. This trade-off is explainably illustrated in the visualization, which only requires testing on batches of test data. In contrast, conventional single-policy methods require entire retraining to incorporate new objectives and rely on extensive trial-and-error to determine each objective’s importance.
@@
EMORL Delivers Comparable Performance in Performance Metrics.
@@
On the PAIR dataset, EMORL achieves an average score of $0.7907$, outperforming parameter-level aggregation Model Soups ($0.6982$) and performing comparably to the single-policy Uniform Weighted ($0.8397$) and DynaOpt ($0.8254$) methods. Notably, model performance varies slightly between the two datasets. On Psych8k, EMORL achieves the mean reward of $0.8082$, with its reflection score reaching the highest ($0.9784$) among all models. Compared to Model Soups, EMORL shows a substantial improvement of $0.25$ in fluency scores, demonstrating the effectiveness of hidden-state aggregation.
@@
The sample generations are shown in Figure 6. EMORL improves reflection and empathy by employing second-person speech ("you"), introducing new perspectives ("job", "office"), and crafting understandable and empathetic statements ("feeling") in response to prompts. EMORL can paraphrase prompts by reflecting on the seeker’s problem and aligning well with the desired objectives for generations, highlighting its effectiveness.
7 Discussion and Conclusion
To conclude, our study addresses the challenges of multi-objective optimization in LLM fine-tuning. We identify the key limitations of convergence speed, process stability, and performance metrics in conventional single-policy approaches, where multiple linguistic objectives are combined into one reward function. To address these limitations, we propose EMORL, a novel meta-policy framework using ensemble learning to aggregate diverse models trained on individual objectives. The results demonstrate that EMORL achieves greater diversity, efficiency, scalability, and explainability while maintaining performance comparable to existing methods in counselor generation tasks.
Our approach is the first to aggregate hidden states to incorporate multiple objectives in NLP. Unlike classification or regression models, which can simply concatenate results from each model to collaboratively decide on the final output, LLMs produce complete sequential generations where fluency cannot be separated or neglected. The hidden states represent rich contextual information, which proves more valuable than incorporating parameters or logits. By using the hidden-state level aggregation, EMORL exceeds the performance limitations of other meta-policy methods. EMORL also offers unique advantages over conventional single-policy methods, including improved training efficiency, extensive scalability, and enhanced explainability due to the parallel training properties, making it more flexible and interpretable than conventional single-policy methods. We investigated the nature of the multi-objective optimization problem and discovered that optima consistently appear within higher-performance regions. This insight led us to design a simple yet effective hierarchical grid search algorithm that requires fewer evaluations to find the globally optimal weights.
Our approach is both task- and model-agnostic. It’s compatible with all transformer-based LLMs, as these architectures maintain hidden states that represent contextual information encompassing multiple objectives in NLP. This study demonstrates the potential of ensemble learning to advance current RL training paradigms and points to a promising novel direction for efficient and flexible multi-objective optimization in future applications.
@@ 7 Discussion and Conclusion
We suggest that our models are not advocated for deployment in clinical or mental health settings. This is because human understanding and communication are indispensable in these domains, and the behavior of language models remains incompletely explored. Instead, we propose that our method and models be utilized for methodological research.
@@ Acknowledgements
We would like to extend our sincere gratitude to BSc. Yaoting Zhu from the Mental Health Faculty at Zhejiang Sci-Tech University for her valuable support on human evaluation, which greatly contributed to the development of this work.
@@ Appendix A Appendix
Table 4: Demonstration of EMORL’s best-performing weight combinations for 5 model pairs in the experiments, along with their average combination.
@@
Table 5: PAIR and Psych8k datasets statistics.
statistics PAIR dataset Psych8k dataset
# of Exchange Pairs 2,544 8,187
Avg # of Words 32.39 45.18
@@ A.1 Parameter-level aggregation
Model Soups, a prominent ensemble learning method, utilizes this parameter-level aggregation strategy for optimizing hyperparameter configurations. However, when applied to multi-objective optimization, this approach yielded suboptimal results. As illustrated in Figure 7, the overall mean reward achieved only $0.6982$, significantly lower than our hidden-state level aggregation method. The fluency metric performed particularly poorly, reaching merely $0.4324$. This underperformance likely stems from the fundamental incompatibility between diverse model objectives at the parameter-level, ultimately failing to effectively combine multiple optimization objectives simultaneously.
Figure 7: Parameter-level aggregation results.
Figure 8: Logit-level aggregation results.
@@ A.2 Logit-level aggregation
We further explored logit-level aggregation as an alternative approach, which is grounded in the work of Shi et al. (2024). This method aggregates the logits, which represent token probabilities across the vocabulary and directly influence token generation. As illustrated in Figure 8, this approach performed even worse for combining multiple objectives, achieving a maximum mean reward of only $0.5934$, with the fluency metric scoring a mere $0.1575$. This poor performance can be attributed to the naive combination of vocabulary probabilities, where predicted token distributions from different local models are simply merged. This process severely impacts fluency, as the resulting text lacks coherence when calculated from disconnected probability distributions. In contrast, our last hidden states aggregation proves more effective by preserving and combining high-level contextual representations during generation, maintaining semantic consistency while still incorporating multiple objectives.
@@ A.4 Evaluation Instruction
The human evaluation is supported by two annotators, one is from China, and the other is from Germany. The evaluation, based on their cross-cultural understanding, supports the robust human-annotated results. This evaluation instruction is derived from the original instruction presented in the work Pérez-Rosas et al. (2024). When evaluating responses, choose the most appropriate score (0, 1, or 2) based on these criteria. Responses may vary in complexity, and the judgment should be guided by the degree to which they reflect upon the client’s prompt.

### v3

@@ 5.1 Models
To evaluate EMORL, we employ T5-base [^1] https://huggingface.co/google-t5/t5-base (220M parameters, Encoder-Decoder architecture) as the pre-trained model. Min et al. (2024) demonstrated T5’s effectiveness for understanding and generating text in counseling tasks, making it suitable for our experiments compared to existing baselines. We utilize Self-Critical Sequence Training (SCST) as the RL algorithm, which generates candidate outputs and computes their mean reward as a baseline, thereby encouraging outputs to exceed this baseline performance (Laban et al., 2021). KL-divergence is incorporated in the loss function to constrain the fine-tuned model from diverging too far from the reference pre-trained model.
In the experiments, we fine-tune models using the default average utility function for reflection, empathy, and fluency across $5$ random seeds, with $3$ generation runs, a training batch size of $16$, and up to $10,000$ steps while implementing early stopping at model convergence. We also employ LoRA Hu et al. (2022) to efficiently manage parameter updates by representing them via low-rank matrices and a scaling factor, as illustrated in Equation (5). Thus, we load the pre-trained model only once during inference and apply the LoRA parameters to update the pre-trained model toward each objective, significantly reducing the computational burden.
For evaluation, we compared EMORL models against four baselines: T5-base, Uniform Weighted, DynaOpt (Min et al., 2024), and Model Soups (Wortsman et al., 2022). All experiments were conducted on a Tesla V100 GPU with 32GB memory, 8 CPU cores, and 40GB system memory, with detailed consumption reported in Table 1.
@@ 5.2 Datasets
Counselor reflection generation task is a single-turn task to generate therapist-like reflective responses that accurately paraphrase and validate a client’s expressed thoughts and feelings (O’neil et al., 2023). The responses are expected to be reflective, empathic to the problem, and maintain fluency, as depicted in Figure 1. These generation objectives frequently conflict with one another. For instance, prioritizing accurate reflection of patient concerns makes maintaining empathetic tone challenging when addressing distressing truths. Similarly, excessive empathetic language may compromise coherence with other response components, undermining overall fluency and discourse quality.
The PAIR [^2] https://lit.eecs.umich.edu/downloads.html dataset is our primary dataset, split into a ratio of $80\%$, $10\%$, and $10\%$, for fine-tuning, aggregation, and inference, respectively (Min et al., 2022). It contains $2,544$ single-turn client-counselor exchanges, covering topics ranging from mental health to lifestyle concerns like diet, exercise, and personal development. To assess the models robustly, we also conduct evaluations on the Psych8k [^3] https://huggingface.co/datasets/EmoCareAI/Psych8k dataset, sampling $10\%$ of its $8,187$ conversation pairs for inference. This dataset focuses on mental health interactions, including anxiety, depression, relationship issues, and stress management. It is widely used for training and evaluating LLMs in mental health counseling, and we leverage it here for reflection generation. The statistics of the datasets are presented in Table 4 (Section A).
Figure 5: The visualization illustrates the hierarchical grid search process, showing the transition from broad search spaces to more refined spaces, where optimal weight combinations are identified. The red line on the color map indicates the maximum objective reward achieved during the search.
@@ 5.3 Metrics
We evaluate comprehensive metrics beyond performance, focusing on six key aspects: (1) Diversity-2 measures linguistic diversity; (2) Edit rate quantifies the avoidance of verbatim repetition; (3) Data consumption tracks the cumulative number of training samples processed; (4) Time consumption records the wall-clock time for each training iteration; (5) Scalability assesses the model’s ability and flexibility to incorporate new preferences and additional objectives; (6) Explainability examines the interpretability of how each objective contributes to the final output.
For performance metrics, we employ text classification LLMs to score objectives on a scale from $0.0$ to $1.0$: (1) Reflection is assessed by the "roberta-base [^4] https://huggingface.co/FacebookAI/roberta-base" with checkpoints from Min et al. (2024), which evaluates the relevance and contextual appropriateness. (2) Empathy is measured by the "bert-empathy [^5] https://huggingface.co/MoaazZaki/bert-empathy", which gauges emotional resonance and understanding. (3) Fluency is evaluated using "gpt2666https://huggingface.co/openai-community/gpt2" by computing the inverse of perplexity, ensuring linguistic smoothness.
@@ 6 Results and Discussion
Hierarchical Grid Visualization Demonstrates Explainable Results:
Figure 5 illustrates the hierarchical grid search process: in iteration $1$ (left subplot), we evaluated $3\times 3\times 3$ weighted combinations for reflection, empathy, and fluency, with color mapping indicating objective reward values. The $2\times 2\times 2$ area with best total performance was then used to refine the search space for subsequent iterations. The right subplot shows iteration $5$, where the search converged to a more precise and refined space: $(0.75,0.8125)$, $(0.4375,0.5)$ and $(0.0,0.0625)$ for the aggregation weights of reflection, empathy, and fluency models respectively. This progressive refinement allows for the precise identification of the optimal weighted combination.
The visualization reveals significant insights regarding each objective’s contribution to overall performance. Unlike defining weight coefficients in the reward function during training, which doesn’t correspond linearly to performance, our aggregation phase’s contribution distribution accurately reflects each objective-specific model’s actual contribution. As shown in Figure 5, the aggregation benefits (orange points) from higher weights of the reflection model. Empathy delivers optimal overall performance with moderate weights, while extreme weights reduce the objective reward. The fluency model demonstrates a negative effect (cyan points) on other objectives when assigned high weights, with lower weights facilitating better integration with the other objectives. This visualization can be readily acquired during aggregation by evaluating weighted combinations on batches of test data. Although results vary across experiments with different sampled models, as shown in Table 5 (Section A), this approach underscores the interpretability and explainability of EMORL.
EMORL Shows Promise Across Multiple Evaluation Metrics:
$T_{total}=\max_{i\in\{1,2,3\}}\{T_{train}(obj_{i})\}+T_{agg}$ (4)
This parallel process enables faster training with over $5869$ fewer data points compared to single-policy methods. While another meta-policy method, Model Soups, employs gradient-based algorithms (learned soup) that often struggle with local optima, leading to increased consumption. However, EMORL employs hidden-state level aggregation and requires token-by-token generation, which interrupts the sequential generation process of transformers, slightly increasing time consumption. EMORL exhibits greater stability, with variations of only $1,650$ data points and $147.43$ seconds, significantly lower than DynaOpt’s variations of $16,736$ data points and $4,365.64$ seconds. This stability is attributed to EMORL’s consistent single-objective fine-tuning and uniform optimization resource consumption facilitated by hierarchical grid search. These advantages position EMORL as an efficient and stable MOO approach.
EMORL is also a flexible fine-tuning approach that achieves both scalability and explainability. Its scalability is demonstrated through two capabilities: 1) Different preferences: EMORL avoids retraining by optimizing weights in the aggregation phase. While our default uses average preferences, user-specific preferences generate new optimal weights during the aggregation phase without costly fine-tuning. 2) New objectives: When incorporating new objectives like helpfulness, EMORL only trains the new objective-specific model and optimizes weights for combining all models, while preserving existing trained models. For explainability, EMORL provides clear insights through hierarchical grid search visualization, as shown in Figure 5. It also demonstrates how increasing one objective’s weight enhances its performance while affecting others. This trade-off is efficiently interpreted through testing only on small data batches. In contrast, conventional single-policy methods require complete retraining for different preferences or new objectives, plus extensive trial-and-error to determine objective importance.
EMORL Delivers Comparable Performance in Performance Metrics:
@@
On the PAIR dataset, EMORL achieves an average score of $0.7907$, outperforming parameter-level aggregation Model Soups ($0.6982$) and performing comparably to the single-policy Uniform Weighted ($0.8397$) and DynaOpt ($0.8254$) methods. Notably, model performance varies slightly between the two datasets. On Psych8k, EMORL achieves the objective reward of $0.8082$, with its reflection score reaching the highest ($0.9784$) among all models. Compared to Model Soups, EMORL shows a substantial improvement of $0.25$ in fluency scores, demonstrating the effectiveness of hidden-state level aggregation for LLM fine-tuning.
@@
The sample generations are shown in Figure 6. EMORL improves reflection and empathy by employing second-person speech ("you"), introducing new perspectives ("job", "office"), and crafting understandable and empathetic statements ("feeling") in response to prompts. EMORL can paraphrase prompts by reflecting on the patient’s problem and aligning well with the desired objectives for generations, highlighting its effectiveness.
7 Conclusion
To conclude, our study addresses the challenges of MOO in LLM fine-tuning. We identify the key limitations of convergence speed, process stability, and performance metrics in conventional single-policy approaches. To address these limitations, we propose EMORL, a novel meta-policy framework using ensemble learning to aggregate diverse models trained on individual objectives. The results demonstrate that EMORL achieves greater diversity, efficiency, scalability, and explainability while maintaining performance comparable to existing methods in counselor generation tasks.
Our approach is the first to aggregate hidden states to incorporate multiple objectives in NLP. Unlike classification or regression models, which can simply concatenate results from each model to collaboratively decide on the final output, LLMs produce complete sequential generations where fluency cannot be separated or neglected. The hidden states represent rich contextual information, which proves more valuable than incorporating parameters or logits. By using the hidden-state level aggregation, EMORL exceeds the performance limitations of meta-policy methods. EMORL also offers unique advantages over conventional single-policy methods, including improved training efficiency, extensive scalability, and enhanced explainability due to the parallel training properties, making it more efficient and flexible. We investigated the nature of the MOO problem and discovered that optima consistently appear within higher-performance regions. This insight led us to design a simple yet effective hierarchical grid search algorithm that requires fewer evaluations to find the globally optimal weights.
Our approach is both task- and model-agnostic. It’s compatible with all transformer-based LLMs, as these architectures maintain hidden states that represent contextual information encompassing multiple objectives in NLP. This study demonstrates the potential of ensemble learning to advance current RL training paradigms and points to a promising novel direction for efficient and flexible MOO in future applications.
@@ 7 Conclusion
We suggest that our models are not advocated for deployment in clinical or mental health settings. This is because human understanding and communication are indispensable in these domains, and the behavior of language models remains incompletely explored. Instead, we propose that our method and models be utilized for methodological research. In future real-world applications, we recommend making users explicitly aware that they are communicating with AI-generated content that does not represent genuine human intelligence or empathy, thereby mitigating ELIZA effects where users may incorrectly attribute human-like qualities to AI systems.
@@ Acknowledgements
This work has been supported by the ARGUS EU project (Grant Agreement No. 101132308), funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or of the European Research Executive Agency (REA). Neither the European Union nor the granting authority can be held responsible for them. The contribution of Cong Yang was supported by the National Natural Science Foundation of China (Grant No. 62473276), the Natural Science Foundation of Jiangsu Province (Grant No. BK20241918), and the Research Fund of Horizon Robotics (Grant No. H230666). We thank Yaoting Zhu from Zhejiang Sci-Tech University for her support on human evaluation.
@@ Appendix A Appendix
Table 4: PAIR and Psych8k datasets statistics.
statistics PAIR dataset Psych8k dataset
# of Exchange Pairs 2,544 8,187
Avg # of Words 32.39 45.18
Table 5: Demonstration of EMORL’s optimal weight combinations for 5 model pairs with different seeds in the experiments, along with their average combination.
@@
Figure 7: Parameter-level aggregation results.
Figure 8: Logit-level aggregation results.
@@ A.1 Parameter-level aggregation
Model Soups (Wortsman et al., 2022), a prominent ensemble learning method, utilizes this parameter-level aggregation strategy for optimizing hyperparameter configurations. However, when applied to MOO, this approach yielded suboptimal results. As illustrated in Figure 7, the overall objective reward achieved only $0.6982$, significantly lower than our hidden-state level aggregation method. The fluency metric performed particularly poorly, reaching merely $0.4324$. This underperformance likely stems from the fundamental incompatibility between diverse model objectives at the parameter-level, ultimately failing to effectively combine multiple optimization objectives.
@@ A.2 Logit-level aggregation
We further explored logit-level aggregation as an alternative approach, which is grounded in the work of Shi et al. (2024). This method aggregates the logits, which represent token probabilities across the vocabulary and directly influence token generation. As illustrated in Figure 8, this approach performed even worse for combining multiple objectives, achieving a maximum objective reward of only $0.5934$, with the fluency metric scoring a mere $0.1575$. This poor performance can be attributed to the naive combination of vocabulary probabilities, where predicted token distributions from different local models are simply merged. This process severely impacts fluency, as the resulting text lacks coherence when calculated from disconnected probability distributions. In contrast, our hidden states aggregation proves more effective by preserving and combining high-level contextual representations during generation, maintaining semantic consistency while still incorporating multiple objectives.
@@ A.4 Evaluation Instruction
The human evaluation is supported by two annotators, one is from China, and the other is from Germany. The evaluation, based on their cross-cultural understanding, supports the robust human-annotated results. This evaluation instruction is derived from the original instruction presented in the work Min et al. (2024). When evaluating responses, choose the most appropriate score (0, 1, or 2) based on these criteria. Responses may vary in complexity, and the judgment should be guided by the degree to which they reflect upon the client’s prompt.